OpenStudy (diyadiya):

Let A=[{0 1}{0 0}] show that (aI + bA)^n =a^nI+na^n-1bA Where I is the identity matrix of order 2

OpenStudy (diyadiya):

\[A=\left[\begin{matrix}0 & 1 \\ 0 & 0\end{matrix}\right]\] Show that \( (aI + bA)^n = a^n I + na^{n-1} bA \)

OpenStudy (amistre64):

of order 2? and as far as I can tell, an exponent on a matrix would imply a diagonalization of eugene values .... maybe

OpenStudy (diyadiya):

Yeah 2x2

OpenStudy (diyadiya):

and i didnt understand the last part

OpenStudy (amistre64):

\begin{pmatrix}a&0\\0&a\end{pmatrix}+\begin{pmatrix}0&b\\0&0\end{pmatrix}

OpenStudy (amistre64):

that didnt format like I wanted :)

OpenStudy (diyadiya):

Lol

OpenStudy (amistre64):

\begin{pmatrix}a&b\\0&a\end{pmatrix}^n is what I hope I see correctly so far

OpenStudy (diyadiya):

I think i have to prove it using Mathematical Induction

OpenStudy (amistre64):

ugh! thats a bit above my abilities at the moment; just starting the linearA stuff this term.

OpenStudy (diyadiya):

Oh :(

OpenStudy (amistre64):

show that its true for some n; then work it out for n+1 is what I can recall from math induction

OpenStudy (diyadiya):

Yeah ,For n then k & k+1

OpenStudy (amistre64):

its the extra variables, a and b thatll throw me for a loop

OpenStudy (amistre64):

good luck with it :)

OpenStudy (diyadiya):

Haha , :D Thanks !!

OpenStudy (anonymous):

does it mean this \[\begin{pmatrix}a&b\\0&a\end{pmatrix}^n=\begin{pmatrix}a^n&na^{n-1}b\\0&a^n\end{pmatrix}\]??

OpenStudy (anonymous):

\[\begin{pmatrix}a&b\\0&a\end{pmatrix}\times \begin{pmatrix}a&b\\0&a\end{pmatrix}=\begin{pmatrix}a^2&2ab\\0&a^2\end{pmatrix}\] by computation, so it seems good

OpenStudy (diyadiya):

yeah, i get that when i substitute the matrix in (aI + bA)^n =a^nI+na^n-1bA

OpenStudy (anonymous):

you could also use this step as the case n = 2 to start your induction proof, or you could use n = 1 which is trivial, too trivial really, so it will be a proof by induction right?

OpenStudy (anonymous):

say we know it is true for all \[k\leq n\] and we want to prove it is true for \[k=n+1\]

OpenStudy (anonymous):

then write \[\begin{pmatrix}a&b\\0&a\end{pmatrix}^{n+1}=\begin{pmatrix}a&b\\0&a\end{pmatrix}^n\times \begin{pmatrix}a&b\\0&a\end{pmatrix}\]

OpenStudy (anonymous):

"by induction" we know it is true for \[\begin{pmatrix}a&b\\0&a\end{pmatrix}^n\] i.e. we know that \[\begin{pmatrix}a&b\\0&a\end{pmatrix}=\begin{pmatrix}a^n&na^{n-1}b\\0&a^n\end{pmatrix}\]

OpenStudy (anonymous):

so finish by multiplying \[\begin{pmatrix}a&b\\0&a\end{pmatrix}^{n+1}=\begin{pmatrix}a&b\\0&a\end{pmatrix}^n\times \begin{pmatrix}a&b\\0&a\end{pmatrix}\] \[=\begin{pmatrix}a^n&na^{n-1}b\\0&a^n\end{pmatrix}\times \begin{pmatrix}a&b\\0&a\end{pmatrix}\] and the last step is matrix multiplication, which should give you the necessary answer

OpenStudy (anonymous):

namely \[=\begin{pmatrix}a^n&na^{n-1}b\\0&a^n\end{pmatrix}\times \begin{pmatrix}a&b\\0&a\end{pmatrix}= \begin{pmatrix}a^{n+1}&(n+1)a^nb\\0&a^{n+1}\end{pmatrix}\]

OpenStudy (anonymous):

which it does!

OpenStudy (diyadiya):

Oh Alright :D Thankyou so Much :)

OpenStudy (anonymous):

yw, hope it is clear.

OpenStudy (diyadiya):

Yes Very clear :)